# A Unified Photometric Redshift Calibration for Weak Lensing Surveys using the Dark Energy Spectroscopic Instrument

### v1

@@ 3.1.2 Importance Weights
In the above equation, $0\leq p_{\mathrm{f,min}}\leq 1$ is a free variable. In essence, targets with a fiber fraction above $p_{\mathrm{f,min}}$ only enter our sample if they received a fiber by DESI and, if they do, are upweighted by $p_{\mathrm{f}}^{-1}$. We note that this includes galaxies that were observed spectroscopically by DESI but did not obtain a successful, i.e., high-confidence redshift measurement. The rationale for this decision is that the DESI instrument is optimized for certain galaxies, e.g., luminous red galaxies or emission-line galaxies, at $z\lesssim 1.6$ (DESI Collaboration et al., 2016a). Only using objects with high-confidence DESI redshifts would underestimate the number of galaxies at $z\gtrsim 1.6$ (Blanco et al., in prep.) and would likely lead to biases. This is issue is not limited to DESI as only using high-confidence spectroscopic redshifts from other surveys can lead to biases in weak lensing calibrations, as well (Hartley et al., 2020). Finally, objects with a low fiber fraction enter our calibration sample regardless of whether they were observed by DESI or not, i.e., this includes objects that only have COSMOS2020 multi-band photometric redshifts. Those objects are not upweighted. Adjusting $p_{\mathrm{f,min}}$ changes how much of our sample is calibrated by DESI. Small values of $p_{\mathrm{f,min}}$ correspond to a large fraction of the sample having DESI spectra at the expense of larger shot noise and potential biases. Large values $p_{\mathrm{f,min}}$ will correspond to a lower fraction of the weight coming from galaxies observed with DESI while having a lower shot noise. In particular, $p_{\mathrm{f,min}}=100\%$ corresponds to taking the entire COSMOS2020 catalog without importance weighting, i.e., $w_{\mathrm{c}}\equiv 1$.
3.2 Lensing Weights
@@
    Not every galaxy in the COSMOS field is expected to be part of a weak lensing sample $b$. To account for this, we weigh every galaxy by the probability $w_{\mathrm{t}}(b)=p(b)$, the probability for each galaxy to make the weak lensing quality cut and be assigned the tomographic bin associated with $b$. This is akin the so-called transfer function (Myles et al., 2021).
@@
For every galaxy in the COSMOS field, the DES Balrog catalog (Everett et al., 2022) contains a large number of artificial injections $n_{\mathrm{inj}}$ throughout the DES survey footprint. We use these injections to estimate $p(b)$, i.e., $n_{b}/n_{\mathrm{inj}}$, where $n_{b}$ is the number of injections that fall inside $b$. Similarly, we use the Balrog catalog to estimate $\langle w_{\mathrm{s}}w_{R}\rangle_{b}$ through a simple Monte-Carlo estimate. For DES, the response $R$ is estimated through the diagonal of the response matrix, $0.5(R_{11}+R_{22})$ (Huff & Mandelbaum, 2017). Unlike for DES, artificial injections of COSMOS galaxies into the weak lensing survey footprint do not exist for HSC. Instead, the COSMOS area has been observed with conditions expected for the wider HSC weak lensing survey. For every COSMOS galaxy, we thus only have a single realization within the wider weak lensing catalog. Thus, $p(b)$ is $1$ if the galaxy passed the lensing cut and is assigned a certain redshift bin and 0, otherwise. Similarly, $\langle w_{\mathrm{s}}w_{R}\rangle_{b}$ becomes $w_{\mathrm{s}}w_{R}$ and since the shear response is not available for HSC, we just use $w_{\mathrm{s}}$. Finally, the weights for KiDS are estimated in a similar fashion as for HSC. However, unlike for HSC, no lensing cuts or weights have been determined for galaxies in the COSMOS field. Thus, we learn the relation between photometry in the the raw KiDS DR4 photometric catalog and the total lensing weight $w_{\mathrm{l}}$ in the KiDS-1000 SOM-gold catalog using neural networks, i.e., with the MLPRegressor class in scikit-learn, a $\tanh$ activation function, early stopping, no L2 regularization, a batch size of $10,000$, an optimization tolerance of 0 and $3$ hidden layers of $128$, $64$, and $32$ neurons, each.
3.3 Redshift Calibration
@@ 3.4 Statistical Uncertainties
We note that our photometric redshift calibration may also be affected by other systematic uncertainties such as the photometric calibration in the calibration fields as well as errors in the redshifts of galaxies in our calibration sample. We will return to these systematic uncertainties in section 6.
3.4.1 Shot Noise
@@
3.4.2 Cosmic Variance
Matter and galaxies are spatially clustered, forming over- and underdensities on large scales. As a result, certain redshifts may be over- or underrepresented in our calibration sample in a way that is independent of shot noise (Moster et al., 2011). To account for this, we first calculate the covariance $C_{ij}$ in the relative matter overdensity between two volumes at redshifts $z_{i}$ and $z_{j}$. The matter variance depends on the matter correlation function $\xi(r,z)$, where $r$ is the comoving distance between two points (Moster et al., 2011),
@@ 3.4.2 Cosmic Variance
In the above equation, $d$ refers to the comoving distance, $\tilde{d}_{ij}=0.5(d_{i}+d_{j})$ and $f(\theta)$ is the normalized distribution of separation angles between two random points inside the survey footprint. The distribution $f(\theta)$ can be directly estimated from the COSMOS calibration catalog assuming that galaxies are only weakly clustered. Using CAMB (Lewis et al., 2000; Howlett et al., 2012) version 1.5.5 to calculate the linear matter power spectrum at different cosmic times, we can then calculate the matter covariance between different redshifts $z_{i}$ and $z_{j}$.
@@ 4 Verification
4.1 Fiber Assignment Probabilities
@@
A key ingredient of our analysis is an accurate estimate of the DESI fiber probability of galaxies in the COSMOS field. As a first step, we perform a visual inspection via SOMs to verify that the results predicted by the deep learning model are plausible. We use SOMs as compared to, e.g., single-band magnitudes or two-band colors, because they allow us to study fiber probabilities in high-dimensional color and magnitude space. [^1] We also studied the grizygrizy magnitude, color, and color-color distributions of DES/KiDS targets matched to COSMOS2020, similar, to Fig. 2. We compared distributions with and without importance weights and found them to be in very good agreement. In Fig. 3, we show the completeness in DESI fiber assignment for galaxies in the DES and KiDS COSMOS catalogs. For DES, galaxies are projected onto the deep DES SOM (Myles et al., 2021) and KiDS galaxies are projected onto the KiDS SOM (Wright et al., 2020). The left-hand panel shows the actual fiber fraction in the SOM whereas the middle panel shows the predicted completeness. The predicted completeness in each SOM cell is simply the average $p_{\mathrm{f}}$ of all galaxies in each cell. It is worth emphasizing that these are pure predictions. As described in section 3.1.1, we used a K-fold approach. As a result, the $p_{\mathrm{f}}$ estimates for each galaxy are never obtained from a model that was trained on the galaxy itself. Overall, the network correctly predicts the trends seen in the data. The right-hand panel, showing the difference between data and model predictions, shows no strong trends with color. To evaluate this agreement more quantitatively, we calculate the coefficient of determination $R^{2}$ for the predicted and observed fiber fractions within each SOM cell. We find $R^{2}$ to be $0.816$ and $0.665$ for DES and KiDS, respectively. From Monte-Carlo simulations under the assumption that the fiber fractions are accurate and all fiber assignments are independent of each other, we find that the expected values are $0.813\pm 0.008$ and $0.691\pm 0.007$, respectively. The agreement is very good for DES but the correlation is slightly weaker than expected for KiDS. This may indicate slight inaccuracies in the estimated fiber probabilities but, based on the tests performed in section 4.3 below, we do not think this has a substantial impact on the inferred redshift distributions.
4.2 Impact of Fiber Fraction Threshold
@@
The fiber fraction threshold $p_{\mathrm{f,min}}$ is a key parameter in our method that effectively determines how strongly the redshift calibration relies on galaxies observed by DESI as opposed to galaxies with high-quality multi-band photometric redshifts from COSMOS2020222A similar concept, Pphot, has been used by Speagle et al. (2019) to study and adjust the reliance of the HSC calibration on many-band photometric redshifts as opposed to spectroscopic redshifts.. In Fig. 4, we study how changes in $p_{\mathrm{f,min}}$ affect the relative contribution of DESI redshifts and shot noise for DES, HSC, and KiDS.
@@
4.3 Redshift Distribution
The main goal of this analysis is an accurate recovery of the effective redshift distribution $n(z)$ of different photometric weak lensing samples. Most photometric redshift calibration studies employ mock data from simulations, such as MICE (Carretero et al., 2015) and Buzzard (DeRose et al., 2019), to help with this task, particularly when it comes to verifying the spectroscopic incompleteness correction. However, in this work, we argue that rigorous tests can also be performed on the data itself. The main idea is that high-quality COSMOS2020 photometric redshifts $z_{\mathrm{C20}}$ exist for most galaxies in the COSMOS field. Thus, for every weak lensing catalog, we can directly measure the $z_{\mathrm{C20}}$ redshift distribution in the COSMOS field and test how well we can recover this distribution from a biased subset of galaxies with high-quality redshifts. While COSMOS2020 redshifts have a certain scatter and catastrophic outliers around the true (spectroscopic) redshift (Weaver et al., 2022), especially for faint targets observed with HSC, the correlation with the true redshift is very strong such that one should expect that biases in the recovered $n(z)$ should show up in $n(z_{\mathrm{C20}})$.
@@ 4.3 Redshift Distribution
Finally, we note that the test performed here is primarily a test related to the recovery of intrinsic redshift distributions from biased samples, similar to most tests with mock catalogs (see, e.g., Wright et al., 2020; Myles et al., 2021). It is worth reiterating that for our fiducial analysis, we use COSMOS2020 redshifts which have scatter around true redshifts and may bias our estimates of $n(z)$. The same is true, but to a lesser extent, for redshifts from DESI. Similarly, we do not estimate potential biases in the photometric calibration in the deep fields (see, e.g., Myles et al., 2021). We leave such investigations to future work.
5 Results
@@
In Fig. 6, we show the main result of our study: estimates of the effective redshift distributions for DES, HSC, and KiDS. We also compare these results against the published $n(z)$ of the three weak lensing survey (Hikage et al., 2019; Hildebrandt et al., 2021; Myles et al., 2021). The results are plotted as $n(a)$, where $a=(1+z)^{-1}$ is the scale factor, instead of $n(z)$ for visual clarity. For convenience, we also list the results for the mean redshift in Table 1, including estimates of shot noise and cosmic variance, as well as the difference compared to the published results. For KiDS-1000, the fiducial results take into account the bias as measured from the MICE mocks (Col. 2, Table 3 in Hildebrandt et al., 2021).
5.1 Dark Energy Survey
Overall, we find good agreement with the results from DES (Myles et al., 2021), both in terms of the mean and distribution of redshifts. The strongest difference in the mean redshift occurs for the third tomographic bin, DES-3, with $\Delta\bar{z}_{\mathrm{fid}}=+0.017$. However, the published DES Y3 results incorporate the effects of blending (MacCrann et al., 2022) as well as additional constraints from clustering redshifts (Gatti et al., 2022). When comparing to the DES Y3 results without these effects, the difference between our and the fiducial results reduces further with $\Delta\bar{z}_{\mathrm{fid}}=[-0.010,-0.007,+0.008,-0.004]$ for the 4 tomographic bins (see Table 2 in Myles et al., 2021).
Furthermore, for DES Y3, we can estimate whether the redshift distribution in the overlap region with the DESI main survey footprint, i.e., at declination $\delta>-10\,\mathrm{deg}$, is significantly different from the distribution in the full DES Y3 footprint. To do so, we recompute the lensing weights, $w_{R}$, using only Balrog injections at $\delta>-10\,\mathrm{deg}$. We find only a very small impact on the mean redshifts for all tomographic bins, i.e., $\Delta\bar{z}<0.01$, suggesting that using the full-footprint redshift distributions in DESI galaxy-galaxy lensing studies does not lead to significant biases in the inferred cosmology.
5.2 Hyper Suprime-Cam
@@
5.3 Kilo-Degree Survey
@@
6 Discussion
6.1 Direct Calibration
In this paper, we have introduced a new technique for the direct calibration of photometric redshifts in weak lensing surveys. Whereas the fiducial studies all employ SOMs (Hikage et al., 2019; Hildebrandt et al., 2021; Myles et al., 2021), our method utilizes neural network regression. Despite these apparent differences, conceptually, both methods share a number of similarities.
The deep-wide SOM approach developed in Buchs et al. (2019) uses two sets of photometric data and associated SOMs [^3] HSC and KiDS do not explicitly use the deep-wide SOM approach and, instead, rely on a single SOM. Such a single SOM approach can be regarded as a special case of the deep-wide SOM method where wide and deep SOM and associated photometry are identical.. First, there are deep-photometry fields with more color information and higher depth than the wide weak lensing survey. These deep photometric fields include the calibration sample with high-quality redshifts. One then groups objects into galaxies occupying the same cell in the deep SOM. Within each SOM cell, the galaxies in the calibration sample are assumed to be representative of all galaxies, including those without high-quality redshifts. Effectively, galaxies in the calibration sample are upweighted by $N_{\mathrm{SOM}}/N_{\mathrm{cal}}$, where $N_{\mathrm{SOM}}$ is the number of galaxies in the SOM cell and $N_{\mathrm{cal}}$ the number of galaxies in the SOM cell that also have high-quality redshifts. The deep SOM thus serves the purpose of alleviating selection biases in the calibration sample and the similarities to the importance weights discussed in section 3.1 should be apparent.

### v2

@@ III.1.2 Importance Weights
In the above equation, $0\leq p_{\mathrm{f,min}}\leq 1$ is a free variable. In essence, targets with a fiber fraction above $p_{\mathrm{f,min}}$ only enter our sample if they received a fiber by DESI and, if they do, are upweighted by $p_{\mathrm{f}}^{-1}$. We note that this includes galaxies that were observed spectroscopically by DESI but did not obtain a successful, i.e., high-confidence redshift measurement. The rationale for this decision is that the DESI instrument is optimized for certain galaxies, e.g., luminous red galaxies or emission-line galaxies, at $z\lesssim 1.6$ (DESICollaboration2016_arXiv_1611_0036). Only using objects with high-confidence DESI redshifts would underestimate the number of galaxies at $z\gtrsim 1.6$ (Blanco et al., in prep.) and would likely lead to biases. This is issue is not limited to DESI as only using high-confidence spectroscopic redshifts from other surveys can lead to biases in weak lensing calibrations, as well (Hartley2020_MNRAS_496_4769). Finally, objects with a low fiber fraction enter our calibration sample regardless of whether they were observed by DESI or not, i.e., this includes objects that only have COSMOS2020 multi-band photometric redshifts. Those objects are not upweighted. Adjusting $p_{\mathrm{f,min}}$ changes how much of our sample is calibrated by DESI. Small values of $p_{\mathrm{f,min}}$ correspond to a large fraction of the sample having DESI spectra at the expense of larger shot noise and potential biases. Large values $p_{\mathrm{f,min}}$ will correspond to a lower fraction of the weight coming from galaxies observed with DESI while having a lower shot noise. In particular, $p_{\mathrm{f,min}}=100\%$ corresponds to taking the entire COSMOS2020 catalog without importance weighting, i.e., $w_{\mathrm{c}}\equiv 1$.
III.2 Lensing Weights
@@
    Not every galaxy in the COSMOS field is expected to be part of a weak lensing sample $b$. To account for this, we weigh every galaxy by the probability $w_{\mathrm{t}}(b)=p(b)$, the probability for each galaxy to make the weak lensing quality cut and be assigned the tomographic bin associated with $b$. This is akin the so-called transfer function (Myles2021_MNRAS_505_4249).
@@
For every galaxy in the COSMOS field, the DES Balrog catalog (Everett2022_ApJS_258_15) contains a large number of artificial injections $n_{\mathrm{inj}}$ throughout the DES survey footprint. We use these injections to estimate $p(b)$, i.e., $n_{b}/n_{\mathrm{inj}}$, where $n_{b}$ is the number of injections that fall inside $b$. Similarly, we use the Balrog catalog to estimate $\langle w_{\mathrm{s}}w_{R}\rangle_{b}$ through a simple Monte-Carlo estimate. For DES, the response $R$ is estimated through the diagonal of the response matrix, $0.5(R_{11}+R_{22})$ (Huff2017_arXiv_1702_2600). Unlike for DES, artificial injections of COSMOS galaxies into the weak lensing survey footprint do not exist for HSC. Instead, the COSMOS area has been observed with conditions expected for the wider HSC weak lensing survey. For every COSMOS galaxy, we thus only have a single realization within the wider weak lensing catalog. Thus, $p(b)$ is $1$ if the galaxy passed the lensing cut and is assigned a certain redshift bin and 0, otherwise. Similarly, $\langle w_{\mathrm{s}}w_{R}\rangle_{b}$ becomes $w_{\mathrm{s}}w_{R}$ and since the shear response is not available for HSC, we just use $w_{\mathrm{s}}$. Finally, the weights for KiDS are estimated in a similar fashion as for HSC. However, unlike for HSC, no lensing cuts or weights have been determined for galaxies in the COSMOS field. Thus, we learn the relation between photometry in the the raw KiDS DR4 photometric catalog and the total lensing weight $w_{\mathrm{l}}$ in the KiDS-1000 SOM-gold catalog using neural networks, i.e., with the MLPRegressor class in scikit-learn, a $\tanh$ activation function, early stopping, no L2 regularization, a batch size of $10,000$, an optimization tolerance of 0 and $3$ hidden layers of $128$, $64$, and $32$ neurons, each.
III.3 Redshift Calibration
@@ III.4 Statistical Uncertainties
We note that our photometric redshift calibration may also be affected by other systematic uncertainties such as the photometric calibration in the calibration fields as well as errors in the redshifts of galaxies in our calibration sample. We will return to these systematic uncertainties in section VI.
III.4.1 Shot Noise
@@
III.4.2 Cosmic Variance
Matter and galaxies are spatially clustered, forming over- and underdensities on large scales. As a result, certain redshifts may be over- or underrepresented in our calibration sample in a way that is independent of shot noise (Moster2011_ApJ_731_113). To account for this, we first calculate the covariance $C_{ij}$ in the relative matter overdensity between two volumes at redshifts $z_{i}$ and $z_{j}$. The matter variance depends on the matter correlation function $\xi(r,z)$, where $r$ is the comoving distance between two points (Moster2011_ApJ_731_113),
@@ III.4.2 Cosmic Variance
In the above equation, $d$ refers to the comoving distance, $\tilde{d}_{ij}=0.5(d_{i}+d_{j})$ and $f(\theta)$ is the normalized distribution of separation angles between two random points inside the survey footprint. The distribution $f(\theta)$ can be directly estimated from the COSMOS calibration catalog assuming that galaxies are only weakly clustered. Using CAMB (Lewis2000_ApJ_538_473; Howlett2012_JCAP_04_027) version 1.5.5 to calculate the linear matter power spectrum at different cosmic times, we can then calculate the matter covariance between different redshifts $z_{i}$ and $z_{j}$.
@@ IV Verification
IV.1 Fiber Assignment Probabilities
@@
A key ingredient of our analysis is an accurate estimate of the DESI fiber probability of galaxies in the COSMOS field. As a first step, we perform a visual inspection via SOMs to verify that the results predicted by the deep learning model are plausible. We use SOMs as compared to, e.g., single-band magnitudes or two-band colors, because they allow us to study fiber probabilities in high-dimensional color and magnitude space. [^1] We also studied the grizygrizy magnitude, color, and color-color distributions of DES/KiDS targets matched to COSMOS2020, similar, to Fig. 2. We compared distributions with and without importance weights and found them to be in very good agreement. In Fig. 3, we show the completeness in DESI fiber assignment for galaxies in the DES and KiDS COSMOS catalogs. For DES, galaxies are projected onto the deep DES SOM (Myles2021_MNRAS_505_4249) and KiDS galaxies are projected onto the KiDS SOM (Wright2020_AA_637_100). The left-hand panel shows the actual fiber fraction in the SOM whereas the middle panel shows the predicted completeness. The predicted completeness in each SOM cell is simply the average $p_{\mathrm{f}}$ of all galaxies in each cell. It is worth emphasizing that these are pure predictions. As described in section III.1.1, we used a K-fold approach. As a result, the $p_{\mathrm{f}}$ estimates for each galaxy are never obtained from a model that was trained on the galaxy itself. Overall, the network correctly predicts the trends seen in the data. The right-hand panel, showing the difference between data and model predictions, shows no strong trends with color. To evaluate this agreement more quantitatively, we calculate the coefficient of determination $R^{2}$ for the predicted and observed fiber fractions within each SOM cell. We find $R^{2}$ to be $0.816$ and $0.665$ for DES and KiDS, respectively. From Monte-Carlo simulations under the assumption that the fiber fractions are accurate and all fiber assignments are independent of each other, we find that the expected values are $0.813\pm 0.008$ and $0.691\pm 0.007$, respectively. The agreement is very good for DES but the correlation is slightly weaker than expected for KiDS. This may indicate slight inaccuracies in the estimated fiber probabilities but, based on the tests performed in section IV.3 below, we do not think this has a substantial impact on the inferred redshift distributions.
IV.2 Impact of Fiber Fraction Threshold
@@
The fiber fraction threshold $p_{\mathrm{f,min}}$ is a key parameter in our method that effectively determines how strongly the redshift calibration relies on galaxies observed by DESI as opposed to galaxies with high-quality multi-band photometric redshifts from COSMOS2020222A similar concept, Pphot, has been used by Speagle2019_MNRAS_490_5658 to study and adjust the reliance of the HSC calibration on many-band photometric redshifts as opposed to spectroscopic redshifts.. In Fig. 4, we study how changes in $p_{\mathrm{f,min}}$ affect the relative contribution of DESI redshifts and shot noise for DES, HSC, and KiDS.
@@
IV.3 Redshift Distribution
The main goal of this analysis is an accurate recovery of the effective redshift distribution $n(z)$ of different photometric weak lensing samples. Most photometric redshift calibration studies employ mock data from simulations, such as MICE (Carretero2015_MNRAS_447_646) and Buzzard (DeRose2019_arXiv_1901_2401), to help with this task, particularly when it comes to verifying the spectroscopic incompleteness correction. However, in this work, we argue that rigorous tests can also be performed on the data itself. The main idea is that high-quality COSMOS2020 photometric redshifts $z_{\mathrm{C20}}$ exist for most galaxies in the COSMOS field. Thus, for every weak lensing catalog, we can directly measure the $z_{\mathrm{C20}}$ redshift distribution in the COSMOS field and test how well we can recover this distribution from a biased subset of galaxies with high-quality redshifts. While COSMOS2020 redshifts have a certain scatter and catastrophic outliers around the true (spectroscopic) redshift (Weaver2022_ApJS_258_11), especially for faint targets observed with HSC, the correlation with the true redshift is very strong such that one should expect that biases in the recovered $n(z)$ should show up in $n(z_{\mathrm{C20}})$.
@@ IV.3 Redshift Distribution
Finally, we note that the test performed here is primarily a test related to the recovery of intrinsic redshift distributions from biased samples, similar to most tests with mock catalogs (see, e.g., Wright2020_AA_637_100; Myles2021_MNRAS_505_4249). It is worth reiterating that for our fiducial analysis, we use COSMOS2020 redshifts which have scatter around true redshifts and may bias our estimates of $n(z)$. The same is true, but to a lesser extent, for redshifts from DESI. Similarly, we do not estimate potential biases in the photometric calibration in the deep fields (see, e.g., Myles2021_MNRAS_505_4249). We leave such investigations to future work.
V Results
@@
In Fig. 6, we show the main result of our study: estimates of the effective redshift distributions for DES, HSC, and KiDS. We also compare these results against the published $n(z)$ of the three weak lensing survey (Hikage2019_PASJ_71_43; Hildebrandt2021_AA_647_124; Myles2021_MNRAS_505_4249). The results are plotted as $n(a)$, where $a=(1+z)^{-1}$ is the scale factor, instead of $n(z)$ for visual clarity. For convenience, we also list the results for the mean redshift in Table 1, including estimates of shot noise and cosmic variance, as well as the difference compared to the published results. For KiDS-1000, the fiducial results take into account the bias as measured from the MICE mocks (Col. 2, Table 3 in Hildebrandt2021_AA_647_124).
V.1 Dark Energy Survey
Overall, we find good agreement with the results from DES (Myles2021_MNRAS_505_4249), both in terms of the mean and distribution of redshifts. The strongest difference in the mean redshift occurs for the third tomographic bin, DES-3, with $\Delta\bar{z}_{\mathrm{fid}}=+0.017$. However, the published DES Y3 results incorporate the effects of blending (MacCrann2022_MNRAS_509_3371) as well as additional constraints from clustering redshifts (Gatti2022_MNRAS_510_1223). When comparing to the DES Y3 results without these effects, the difference between our and the fiducial results reduces further with $\Delta\bar{z}_{\mathrm{fid}}=[-0.010,-0.007,+0.008,-0.004]$ for the 4 tomographic bins (see Table 2 in Myles2021_MNRAS_505_4249).
Furthermore, for DES Y3, we can estimate whether the redshift distribution in the overlap region with the DESI main survey footprint, i.e., at declination $\delta>-10\,\mathrm{deg}$, is significantly different from the distribution in the full DES Y3 footprint. To do so, we recompute the lensing weights, $w_{R}$, using only Balrog injections at $\delta>-10\,\mathrm{deg}$. We find shift towards slightly lower mean redshifts, i.e., $\Delta\bar{=}-0.002\pm 0.003$, $008\pm 0.003, 0.0$$009\pm 0.003$, and $0.009\pm 0.005$ for the four tomographic bins. This suggest that using the full-footprint redshift distributions in DESI galaxy-galaxy lensing studies does not lead to significant biases in the inferred cosmology.
V.2 Hyper Suprime-Cam
@@
V.3 Kilo-Degree Survey
@@
VI Discussion
VI.1 Direct Calibration
In this paper, we have introduced a new technique for the direct calibration of photometric redshifts in weak lensing surveys. Whereas the fiducial studies all employ SOMs (Hikage2019_PASJ_71_43; Hildebrandt2021_AA_647_124; Myles2021_MNRAS_505_4249), our method utilizes neural network regression. Despite these apparent differences, conceptually, both methods share a number of similarities.
The deep-wide SOM approach developed in Buchs2019_MNRAS_489_820 uses two sets of photometric data and associated SOMs [^3] HSC and KiDS do not explicitly use the deep-wide SOM approach and, instead, rely on a single SOM. Such a single SOM approach can be regarded as a special case of the deep-wide SOM method where wide and deep SOM and associated photometry are identical.. First, there are deep-photometry fields with more color information and higher depth than the wide weak lensing survey. These deep photometric fields include the calibration sample with high-quality redshifts. One then groups objects into galaxies occupying the same cell in the deep SOM. Within each SOM cell, the galaxies in the calibration sample are assumed to be representative of all galaxies, including those without high-quality redshifts. Effectively, galaxies in the calibration sample are upweighted by $N_{\mathrm{SOM}}/N_{\mathrm{cal}}$, where $N_{\mathrm{SOM}}$ is the number of galaxies in the SOM cell and $N_{\mathrm{cal}}$ the number of galaxies in the SOM cell that also have high-quality redshifts. The deep SOM thus serves the purpose of alleviating selection biases in the calibration sample and the similarities to the importance weights discussed in section III.1 should be apparent.